Q-7An electron is moving at 2 meters per second in the +z direction when it encounters a 12-N/C electric field directed in thquestion 7
Answer #1

velocity of erecson V = amis (R) Electric field - 12 Nic (1) Magnesic field B = 12+ (3) charse on electron, a ne = -1.6x10reMass of election, me - 9.11x10 3 ks @ Electric force Ę - % = -1.641019 (121) = 1.924 18 18 (1 magnitude & 1.924 2018 N NegatiNet force Fier - F + 21.92 +10 18 (-1) + 3.8441018 (9) Frej = 1.9241018 N(7) Magnitude of had force Fret) = 1.9241018 ) accel